# From Freshness to Effectiveness: Goal-Oriented Sampling for Remote Decision Making

### v2

@@ IV-B A Reformulation of Problem 2
The proof is detailed in the the short version of this work in [1, Appendix B]. ∎
@@ 9            $\lambda_{\uparrow}=\lambda$;
IV-C Two-layer Approaches: Bisec-RVI and Bisec-$\tau$-RVI
Following Lemma 2, solving Problem 3 reduces to solving Problem 4 to determine the optimal value $U(\lambda)$ while simultaneously finding the unique root $\rho^{\star}$ of the implicit function $U(\lambda)$, which can be solved using a two-layer nested algorithm as shown in Fig. 3 and Algorithm 1. In the inner layer, value iteration approaches are applied to approximate the optimal value of Problem 4, $U(\lambda)$, by resolving the reformulated MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ detailed in subsection IV-C1. In the outer layer, a bisection search algorithm approximates the unique root $\rho^{\star}$. Conventionally, the RVI algorithm is applied in the inner layer to iteratively approximate $U(\lambda)$ [1, 20], but it proves to be divergent in our formulation (see Fig. 2). To address this limitation, we propose $\tau$-RVI in this paper and present a rigorous convergence analysis. The condition and the rationale of the convergence will be discussed in this section.
IV-C1 Inner-Layer MDP Given Dinkelbach Parameter $\lambda$
@@
  • Transition Probability: The transition probability is defined by $\Pr(\mathcal{G}_{i+1}|\mathcal{G}_{i},Z_{i},A_{i})$. We have the transition probability established in (15), whose detailed proof is given in the short version of this work [1, Appendix C]:
    $\Pr(\mathcal{G}_{i}=(s^{\prime},\delta^{\prime},a^{\prime})|\mathcal{G}_{i}=(s,\delta,a),Z_{i},A_{i})$ (15)
@@
    where $[\mathbf{P}]_{s\times s^{\prime}}$ denotes the element located at the $s$-th row and $s^{\prime}$-th column of the matrix $\mathbf{P}$.
@@
    The proof is given in the short version of this work in [1, Appendix D]. ∎
In this paper, we establish that the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ preserves the critical unichain [^7] An MDP is unichain if the resulting Markov chain under any stationary policy is an aperiodic unichain. property under random communication delays and age-aware system configurations. While the primal MDP, defined by a finite state space $\mathcal{S}$, action space $\mathcal{A}$, and transition probabilities $\mathbf{P}_{a}$, is an aperiodic unichain, it remains unresolved whether this property transfers to $\mathscr{P}_{\text{MDP}}(\lambda)$. Direct verification presents significant computational overhead, as determining the unichain property is proven to be NP-hard [91]. To overcome this challenge, the following theorem demonstrates that the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ inherits the unichain property from the primal MDP, thereby guaranteeing the existence of an optimal stationary deterministic policy.
Theorem 1.
If the primal MDP, characterized by a finite state space $\mathcal{S}$, a finite action space $\mathcal{A}$, and the transition probability $\mathbf{P}_{a}$ is an aperiodic unichain, then the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ is unichain, and an optimal stationary deterministic policy [^8] A policy is deemed stationary when it is independent of the time slot $t$. A policy is deemed deterministic if it a deterministic mapping from the state space to the action space. When an optimal stationary policy exists, solving the policy $\boldsymbol{\psi}^{\lambda}$ is equivalent to solving simply a stationary policy $\phi^{\lambda}$. exists for $\mathscr{P}_{\text{MDP}}(\lambda)$.
@@
The proof is given in the short version of this work in [1, Appendix E]. ∎
For a unichain MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ in which an optimal stationary deterministic policy exists, one can establish the following Average Cost Optimality Equations (ACOE) [92, Eq. 4.1]:
@@
where $V^{\star}(\gamma;\lambda)\in\mathbb{R}$ is the optimal value function and $U(\lambda)\in\mathbb{R}$ is the optimal long-term average value of $\mathscr{P}_{\text{MDP}}(\lambda)$ in (14). By solving $U(\lambda)$ and $V^{\star}(\gamma;\lambda)$ for $\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ from the ACOE (IV-C1), we can obtain the optimal stationary deterministic sampling and remote decision making policy for $\mathscr{P}_{\text{MDP}}(\lambda)$, defined as: $\phi^{*}_{\lambda}:\mathcal{S}\times\mathcal{Y}\times\mathcal{A}\rightarrow\mathcal{Z}\times\mathcal{A}$.
@@
The ACOE in (IV-C1) represents a series of nonlinear equations, which is mathematically intractable to solve explicitly. One can resort to the typical Dynamic Programming (DP)-like RVI algorithm [93] to iteratively generate the sequences $\{U_{K}(\lambda)\}^{K\in\mathbb{N}^{+}}$ and $\{V_{K}(\gamma;\lambda)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}^{K\in\mathbb{N}^{+}}$ that conditionally converge asymptotically to the solutions of the ACOE.
@@
(Restatement of [94, Proposition 4.3.2]). If an MDP satisfies the following conditions:
@@
then the sequences $\{U_{K}(\lambda)\}_{K\in\mathbb{N}}$ and $\{V_{K}(\gamma;\lambda)\}_{K\in\mathbb{N}}$ will converge to the solution to the ACOE (IV-C1):
@@
In the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$, condition (a) in Lemma 4 is ensured by Theorem 1. Nevertheless, the following counterexample demonstrates that condition (b) does not necessarily hold in $\mathscr{P}_{\text{MDP}}(\lambda)$. Specifically, even if the primal MDP characterized by $\left\langle\mathcal{S},\mathcal{A},\mathbf{P}_{a},\mathcal{C}\right\rangle$ is aperiodic, the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ may still exhibit periodicity, which will cause the RVI algorithm to diverge.
@@
(A divergence example of RVI). Consider the parameter setup described in Appendix LABEL:appendixh where the delay is constant with $p=0$. In this case, the optimal policy for the transformed MDP $\mathscr{P}_{\text{MDP}}(10)$ is:
@@
Lemma 4 and Example 1 show that the RVI algorithm [93] may not asymptotically converge to $U(\lambda)$, and the reason is the periodic nature inherent in the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$. Consequently, the existing two-layer Bisec-RVI Algorithm (e.g., [1, Algorithm 1], [71, Algorithm 1], and [84, Algorithm 1]) cannot reliably determine the root $\rho^{\star}$. To circumvent this problem, we propose a new iterative approach, namely $\tau$-RVI, in Iteration 2. This approach eliminates the need for condition (b) in Lemma 4 but guarantees rigorous convergence.
@@
$\tilde{U}_{K+1}(\lambda)$ $=\min_{A_{i},Z_{i}}\bigg{\{}\,g(\gamma^{\text{r}},Z_{i},A_{i};\lambda)+\tau\,\mathbb{E}\big{[}\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma^{\text{r}},Z_{i},A_{i}\big{]}\bigg{\}},$ (24a)
$\tilde{V}_{K+1}(\gamma;\lambda)$ $=(1-\tau)\,\tilde{V}_{K}(\gamma;\lambda)+\min_{A_{i},Z_{i}}\bigg{\{}\,g(\gamma,Z_{i},A_{i};\lambda)+\tau\,\mathbb{E}\big{[}\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma,Z_{i},A_{i}\big{]}\bigg{\}}-\tilde{U}_{K+1}(\lambda),\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},$ (24b)
@@
Fig. 4-(b) illustrates the convergence of $\tau$-RVI across various values of the parameter $\tau$, compared to the standard RVI algorithm [93]. The results demonstrate that $\tau$-RVI overcomes oscillatory behavior encountered by RVI, with convergence rates depending on the selected $\tau$ values. A rigorous convergence analysis of $\tau$-RVI will be presented in in section IV-D.
IV-C2 Outer-Layer Bisection and Bounds on $\rho^{\star}$
@@
The proof is given in the short version of this work in [1, Appendix F]. ∎
IV-D Convergence of $\tau$-RVI
In this subsection, we present convergence results for $\tau$-RVI in Iteration 2. We theoretically show that the generated sequences in $\tau$-RVI will asymptotically approach the solution to the ACOE (IV-C1). To quantify the convergence behavior, we define the relative error for the $K$-th iteration value $\tilde{U}_{K}(\lambda)$ with respect to the ACOE solution $U(\lambda)$ as:
@@ IV-D Convergence of $\tau$-RVI
($i$). For $\forall\lambda\in\mathbb{R}$,
@@
($ii$). For $\forall\lambda\in\mathbb{R}$, $\forall 0<\tau<1$, and $\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$,
@@
which can be solved by the traditional RVI with guaranteed convergence because of its unichain and aperiodic property. Then, comparing (IV-C1) and (32), we can establish the relationship that $\tau\tilde{V}^{\star}(\gamma;\lambda)={V}^{\star}(\gamma;\lambda)$ for $\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ and $\tilde{U}(\lambda)={U}(\lambda)$. See Appendix LABEL:proof:convergence:theorem2 for the proof. ∎
The next theorem characterizes an upper bound on the relative error of the $\tau$-RVI Algorithm, whose proof is included in Appendix LABEL:proof:convergence:theorem2.
@@
where $M$ is a scaling factor, $L$ is defined by (LABEL:Ldefinition). The term $R=\frac{1}{(1-\epsilon)^{1/L}}>1$ captures the asymptotic convergence rate.
@@ V-A Fixed-Point Equations and Iterations
The proof is detailed in the short version of this work in [1, Appendix G]. ∎
@@
[…]
@@
Then, as was done in [1, Algorithm 2], we can conduct the following fixed-point iterations to asymptotically approximate the solution to (34):
@@
Typically, the convergence of fixed-point iterations is established using the Banach Contraction Mapping Theorem [86, Theorem 6]. According to this theorem, if the operator $Q(\cdot):\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}\rightarrow\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}$ is a contraction mapping, i.e., for any $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}$, there exists a constant $10<\alpha<$ 1 such that
$||G(\mathbf{W}_{1})-G(\mathbf{W}_{2})||_{2}\leq\alpha||\mathbf{W}_{1}-\mathbf{W}_{2}||_{2},$ (38)
@@
(Divergence Example of FPBI [1, Algorithm 2]). Consider the parameter setup described in Appendix LABEL:appendixh where the delay is constant with $p=0$. In this case, the fixed-point iteration starting from $\mathbf{W}^{0}=\mathbf{0}$ will diverge, as shown in Fig. 5.
Figure 5: Convergence comparison between FPBI [1, Algorithm 2] and OnePDSI (Iteration 3). FPBI becomes non-expansive and thus diverges. The proposed OnePDSI converges.

### v3

@@ IV-A A Reformulation of Problem 2
See Appendix B.
@@ 9    $\lambda_{\uparrow}=\lambda$;
IV-B Two-layer Approaches: Bisec-RVI and Bisec-$\tau$-RVI
Following Lemma 2, solving Problem 3 reduces to solving Problem 4 to determine the optimal value $U(\lambda)$ while simultaneously finding the unique root $\rho^{\star}$ of the implicit function $U(\lambda)$, which can be solved using a two-layer nested algorithm as shown in Fig. 3 and Algorithm 1. In the inner layer, value iteration approaches are applied to approximate the optimal value of Problem 4, $U(\lambda)$, by resolving the reformulated MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ detailed in subsection IV-B1. In the outer layer, a bisection search algorithm approximates the unique root $\rho^{\star}$. Conventionally, the RVI algorithm is applied in the inner layer to iteratively approximate $U(\lambda)$ [49, 85], but it proves to be divergent in our formulation (see Fig. 2). To address this limitation, we propose $\tau$-RVI in this paper and present a rigorous convergence analysis. The condition and the rationale of the convergence will be discussed in this section.
IV-B1 Inner-Layer MDP Given Dinkelbach Parameter $\lambda$
@@
  • Transition Probability: The transition probability is defined by $\Pr(\mathcal{G}_{i+1}|\mathcal{G}_{i},Z_{i},A_{i})$. We have the transition probability established in (18), whose detailed proof is given in Appendix C:
    $\Pr(\mathcal{G}_{i+1}=(s^{\prime},\delta^{\prime},a^{\prime})|\mathcal{G}_{i}=(s,\delta,a),Z_{i},A_{i})$ (18)
@@
    where $\mathbf{P}]_{s\times s^{\prime}}$ denotes the element located at the $s$-th row and $s^{\prime}$-th column of the matrix $\mathbf{P}$.
@@
    See Appendix D.
In what follows, we refer to the MDP with transition dynamics $\mathbf{P}_{a}$ as the primal MDP, and denote by $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ the transformed (or lifted) MDP. The next theorem provides a simple sufficient condition under which the lifted MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ is unichain [^9] An MDP is said to be unichain if, under any stationary policy, the induced Markov chain has a single recurrent class (with all other states being transient)., thereby ensuring the existence of an optimal stationary deterministic policy.
Theorem 1 (Sufficient condition for a unichain lifted MDP).
Let $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{Y}$ be the finite state, action, and delay sets of the primal MDP. Suppose there exist a state $s^{\star}\in\mathcal{S}$, an integer $m\geq 1$, and a constant $\epsilon>0$ such that, for every initial $(s,a)\in\mathcal{S}\times\mathcal{A}$ and for every admissible length-$m$ sequence $\{A_{t},Z_{t},\delta_{t}\}_{t=0}^{m-1}$:
$\big[\mathbf{P}_{a}^{\delta_{0}}\mathbf{P}_{A_{0}}^{Z_{0}}\cdots\mathbf{P}_{A_{m-2}}^{\delta_{m-1}}\mathbf{P}_{A_{m-1}}^{Z_{m-1}}\big]_{s\times s^{\star}}\geq\ \epsilon.$ (22)
Then, for every stationary deterministic policy $\pi:\mathcal{S}\times\mathcal{Y}\times\mathcal{A}\to\mathcal{A}\times\mathcal{Z}$, the Markov chain on $\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ induced by $\pi$ has a single recurrent class (all other states are at most transient). In particular, the lifted MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ is unichain.
@@
See Appendix E. ∎
If the lifted MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ is a unichain, an optimal stationary deterministic policy exists, and one can establish the following Average Cost Optimality Equations (ACOE) [39, Eq. 4.1]:
@@
where $V^{\star}(\gamma;\lambda)\in\mathbb{R}$ is the optimal value function and $U(\lambda)\in\mathbb{R}$ is the optimal long-term average value of $\mathscr{P}_{\text{MDP}}(\lambda)$ in (17). By solving $U(\lambda)$ and $V^{\star}(\gamma;\lambda)$ for $\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ from the ACOE (IV-B1), we can obtain the optimal stationary deterministic sampling and remote decision-making policy for $\mathscr{P}_{\text{MDP}}(\lambda)$, defined as: $\phi^{*}_{\lambda}:\mathcal{S}\times\mathcal{Y}\times\mathcal{A}\rightarrow\mathcal{Z}\times\mathcal{A}$.
@@
The ACOE in (IV-B1) represents a series of nonlinear equations, which is mathematically intractable to solve explicitly. One can resort to the typical Dynamic Programming (DP)-like RVI algorithm [102] to iteratively generate the sequences $\{U_{K}(\lambda)\}^{K\in\mathbb{N}^{+}}$ and $\{V_{K}(\gamma;\lambda)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}^{K\in\mathbb{N}^{+}}$ that conditionally converge asymptotically to the solutions of the ACOE.
@@
(Restatement of [13, Proposition 4.3.2]). If an MDP satisfies the following conditions:
@@
then the sequences $\{U_{K}(\lambda)\}_{K\in\mathbb{N}}$ and $\{V_{K}(\gamma;\lambda)\}_{K\in\mathbb{N}}$ will converge to the solution to the ACOE (IV-B1):
@@
In the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$, condition (a) in Lemma 4 is ensured by Theorem 1. Nevertheless, the following counter-example demonstrates that condition (b) does not necessarily hold in $\mathscr{P}_{\text{MDP}}(\lambda)$. Specifically, even if the primal MDP characterized by $\left\langle\mathcal{S},\mathcal{A},\mathbf{P}_{a},\mathcal{C}\right\rangle$ is aperiodic, the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ may still exhibit periodicity, which will cause the RVI algorithm to diverge.
@@
(A divergence example of RVI). Consider the parameter setup described in Appendix H where the delay is constant with $p=0$. In this case, the optimal policy for the transformed MDP $\mathscr{P}_{\text{MDP}}(10)$ is:
@@
Lemma 4 and Example 1 show that the RVI algorithm [102] may not asymptotically converge to $U(\lambda)$, and the reason is the periodic nature inherent in the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$. Consequently, the existing two-layer Bisec-RVI Algorithm (e.g., [49, Algorithm 1], [19, Algorithm 1], and [7, Algorithm 1]) cannot reliably determine the root $\rho^{\star}$. To circumvent this problem, we propose a new iterative approach, namely $\tau$-RVI, in Iteration 2. This approach eliminates the need for condition (b) in Lemma 4 but guarantees rigorous convergence.
@@
$\tilde{U}_{K+1}(\lambda)$ $=\min_{A_{i},Z_{i}}\bigg\{\,g(\gamma^{\text{r}},Z_{i},A_{i};\lambda)$
$\quad\quad\quad+\tau\,\mathbb{E}\big[\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma^{\text{r}},Z_{i},A_{i}\big]\bigg\},$ (28a)
$\tilde{V}_{K+1}(\gamma;\lambda)$ $=(1-\tau)\,\tilde{V}_{K}(\gamma;\lambda)$
$\quad+\min_{A_{i},Z_{i}}\bigg\{\,g(\gamma,Z_{i},A_{i};\lambda)$
$\quad+\tau\,\mathbb{E}\big[\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma,Z_{i},A_{i}\big]\bigg\}$
$\quad-\tilde{U}_{K+1}(\lambda),\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},$ (28b)
@@
Fig. 4-(b) illustrates the convergence of $\tau$-RVI across various values of the parameter $\tau$, compared to the standard RVI algorithm [102]. The results demonstrate that $\tau$-RVI overcomes oscillatory behavior encountered by RVI, with convergence rates depending on the selected $\tau$ values. A rigorous convergence analysis of $\tau$-RVI will be presented in in section IV-C.
IV-B2 Outer-Layer Bisection and Bounds on $\rho^{\star}$
@@
See Appendix F.
IV-C Convergence of $\tau$-RVI
In this subsection, we present convergence results for $\tau$-RVI in Iteration 2. We theoretically show that the generated sequences in $\tau$-RVI will asymptotically approach the solution to the ACOE (IV-B1). To quantify the convergence behavior, we define the relative error for the $K$-th iteration value $\tilde{U}_{K}(\lambda)$ with respect to the ACOE solution $U(\lambda)$ as:
@@ IV-C Convergence of $\tau$-RVI
(ii). For $\forall\lambda\in\mathbb{R}$,
@@
(iiii). For $\forall\lambda\in\mathbb{R}$, $\forall 0<\tau<1$, and $\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$,
@@
which can be solved by the traditional RVI with guaranteed convergence because of its unichain and aperiodic property. Then, comparing (IV-B1) and (36), we can establish the relationship that $\tau\tilde{V}^{\star}(\gamma;\lambda)={V}^{\star}(\gamma;\lambda)$ for $\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ and $\tilde{U}(\lambda)={U}(\lambda)$. See Appendix I for the detailed proof. ∎
The next theorem characterizes an upper bound on the relative error of the $\tau$-RVI Algorithm, whose proof is included in Appendix I.
@@
where $M$ is a scaling factor, $L$ is defined by (129). The term $R=\frac{1}{(1-\epsilon)^{1/L}}>1$ captures the asymptotic convergence rate.
@@ V-A Fixed-Point Equations and Iterations
See Appendix G.
@@
$\left\{\begin{aligned} &\mathbf{W}^{\star}=T(\mathbf{W}^{\star},\rho^{\star})\\ &\rho^{\star}=H(\mathbf{W}^{\star}),\end{aligned}\right$. (39)
@@
Then, as was done in [49, Algorithm 2], we can conduct the following fixed-point iterations to asymptotically approximate the solution to (38):
$\left\{\begin{aligned} &\mathbf{W}^{K+1}=G(\mathbf{W}^{K}),\\ &\rho_{K+1}=H(\mathbf{W}^{K+1}).\end{aligned}\right$. (41)
Typically, the convergence of fixed-point iterations is established using the Banach Contraction Mapping Theorem [5, Theorem 6]. According to this theorem, if the operator $Q(\cdot):\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}\rightarrow\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}$ is a contraction mapping, i.e., for any $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}$, there exists a constant $0<\alpha<1$ such that
@@
(Divergence Example of FPBI [49, Algorithm 2]). Consider the parameter setup described in Appendix H where the delay is $p=0$. In this case, the fixed-point iteration starting from $\mathbf{W}^{0}=\mathbf{0}$ will diverge, as shown in Fig. 5.
Figure 5: Convergence comparison between FPBI [49, Algorithm 2] and OnePDSI (Iteration 3). FPBI becomes non-expansive and thus diverges. The proposed OnePDSI converges.